OpenStudy (anonymous):

What is the slope-intercept form of this equation? 8x + 6y = 12

OpenStudy (anonymous):

4x

OpenStudy (alexandervonhumboldt2):

ax+by=12 -a/b=slope

OpenStudy (anonymous):

4x

OpenStudy (alexandervonhumboldt2):

@FaZe_Rainn that incorrect

OpenStudy (diamondboy):

make y feel lonely

OpenStudy (elonasushchik):

1. 8x-6y=12 6y=12 (cover up the "8x") /6 (and divide by 6, giving you) y=2 8x=12 (cover up the "6y") /8 x=1.5 2. -y+2x=6 y=-1 x=3 3. 5y-3x=9 y= -3 x= 1.8 4. y= -2x-5 x=-2.5 y=5

OpenStudy (alexandervonhumboldt2):

@amir7816 that incorrect too

OpenStudy (elonasushchik):

@jayonna999 I explained it to you..

OpenStudy (alexandervonhumboldt2):

@ElonaSushchik @amir7816 @FaZe_Rainn no direct answers

OpenStudy (anonymous):

:(

OpenStudy (elonasushchik):

I explanied it in steps @AlexandervonHumboldt2 im not retarded

OpenStudy (alexandervonhumboldt2):

you should guide no answers

OpenStudy (anonymous):

Soo is it .. A. y = - 4/3x + 2 B. y = -4/3x + 6 C. y = 4/3x + 6 D. y = 2/3x + 2

OpenStudy (anonymous):

@ElonaSushchik

OpenStudy (anonymous):

@ElonaSushchik i hate period

OpenStudy (elonasushchik):

Step 1: Find the slope of the line. To find the slope of the given line we need to get the line into slope-intercept form (y = mx + b), which means we need to solve for y: The slope of the line 8x + 6y = 15 is m = –4/3.

OpenStudy (elonasushchik):

Start with the given equation. I would always put it in order Ax+By=C, but you don't need to. 8x+6y=12 Subtract 8x from both sides. 6y=12-8x Multiply both sides by 1/6 to get y by itself. y=(1/6)(12-8x) Distribute 1/6 y=(1/6)(12)-(1/6)(8)x Multiply y=(12/6)-(8/6)x Rearrange the terms y=(-8/6)x+(12/6) Reduce any fractions y=(-4/3)x+2 In case this looks confusing, where it says (-4/3), that means negative four thirds (or negative four over three) I just don't know how to make fractions on the computer.

OpenStudy (elonasushchik):

@jayonna999 did you understand it?!

OpenStudy (anonymous):

yesss thank you alot , I got it @ElonaSushchik

OpenStudy (elonasushchik):

Your Welcome!!!